# Lightweight Sequential Transformers for Blood Glucose Level Prediction in Type-1 Diabetes

### v1

@@ 3.1 Problem Formulation
3.2 Proposed forecaster model
@@
3.3 Time Sequence Encoder
@@
3.4 Time Sequence Forecaster
@@
it functions identically to the DayTime Embedder described in 3.3. However, instead of processing the daytime of the observed sequence, it handles the daytime corresponding to the element to predict. Specifically, it processes the daytime at time $i+1$ and outputs an embedding $\mathbf{b}_{i+1}$ of dimension $N$.
@@ 4.1 Benchmark datasets
Table 1: Overall performance on OhioT1DM Dataset*
@@ 4.4 Evaluation Metrics
5 Results and Discussion
Table 2: Overall performance on DiaTrend Dataset*
@@ 5 Results and Discussion
Table 1 and Table 2 respectively show the results achieved on the OhioT1DM and DiaTrend datasets, comparing the proposed method and training strategy with the state-of-the-art models. The reported experiments consider both the single and multimodal approaches. All architectures are evaluated on two dataset horizon configurations (30 and 60 minutes). Each table shows the difference in terms of evaluation metrics and parameters, indicating whether the experiment uses only CGM or also additional features. To facilitate the readability of the results, a Ranking column is also reported. The ranking is determined, considering the two PH configurations separately, by the average of the mean metrics’ score, with the score of the parameters. Such scores are computed by normalizing the results of the same metric (same column in the table) between 00 and $1$, considering $1$ as the best result possible and 00 as the worst. The score for the parameters is computed in the same way on the Params column. The average of the scores obtained for 30- and 60-minute horizons are then used to obtain the final rank independently from the horizon considered.
Table 3: Sequential Transformer performance on OhioT1DM and DiaTrend datasets in different training settings*
@@
as observable in Table 1, the best rank is achieved by the proposed model with a single-modality approach, while the multimodal version achieves the third rank. This means that, on the OhioT1DM dataset, the Sequential Transformer proves to be optimal for real-life situations requiring edge computing, thanks to its extremely low dimensionality in terms of parameters and its performance considering only CGM. The second best performing model is Bi-GRU [29]. However, it is important to notice that even if the model is relatively small in terms of parameters and compared with other architectures, its dimension can still be a problem on wearable devices. Another fundamental analysis regards the identification of future adverse events. The proposed strategy with the Sequential Transformer can always achieve the best or among the best results in detecting hyper and hypo events. On this particular issue, it is possible to observe that the multimodal Sequential Transformer outperforms the Bi-GRU, thus being more suitable in a real-life scenario.
@@
Table 2 shows a trend similar to the OhioT1DM dataset. However, in this case, the Sequential Transformer takes both the best and second-best positions on the ranking. The single-modality version achieves the best ranking, with the multimodal version reaching the second place. As before, the Sequential Transformer in every version achieves the best or among the best performance in detecting adverse events. The only exception is TimesNet [27], which achieves the second-best performance in Hypo Sen. Nonetheless, its number of parameters makes it unsuitable for real-life scenarios where edge computing results in a better solution.
5.1 Ablation studies
@@
The first rows of Table 3 report the scores obtained on the OhioT1DM dataset with different settings. The best-ranked method is represented by our Sequential Transformer using the proposed balancing strategy, implementing data augmentation, and, as determined before, using only CGM. The second-best is the multimodal version, with the same setting in terms of loss and data augmentation. It is interesting to notice that, based on the ranking, the balanced versions represent the first four best choices, demonstrating the effectiveness of the defined training strategy. Similar conclusions can be drawn for the DiaTrend dataset. The first rank is still taken by the single-modality setting with a balanced strategy and data augmentation. The second is the same setting but with a multimodal approach. Even in this case, it is possible to notice that the four best-ranked settings are all based on the balance training strategy, remarking once again on the importance of such implementation.
@@
5.2 Clarke Error Grid Analysis
@@
5.3 Deployment on wearable devices

### v2

@@ III-A Problem Formulation
III-B Proposed forecaster model
@@
III-C Time Sequence Encoder
@@
III-D Time Sequence Forecaster
@@
it functions identically to the DayTime Embedder described in III-C. However, instead of processing the daytime of the observed sequence, it handles the daytime corresponding to the element to predict. Specifically, it processes the daytime at time $i+1$ and outputs an embedding $\mathbf{b}_{i+1}$ of dimension $N$.
@@ IV-A Benchmark datasets
TABLE I: Overall performance on OhioT1DM Dataset*
@@ IV-D Evaluation Metrics
V Results and Discussion
TABLE II: Overall performance on DiaTrend Dataset*
@@ V Results and Discussion
Table I and Table II respectively show the results achieved on the OhioT1DM and DiaTrend datasets, comparing the proposed method and training strategy with the state-of-the-art models. The reported experiments consider both the single and multimodal approaches. All architectures are evaluated on two dataset horizon configurations (30 and 60 minutes). Each table shows the difference in terms of evaluation metrics and parameters, indicating whether the experiment uses only CGM or also additional features. To facilitate the readability of the results, a Ranking column is also reported. The ranking is determined, considering the two PH configurations separately, by the average of the mean metrics’ score, with the score of the parameters. Such scores are computed by normalizing the results of the same metric (same column in the table) between 00 and $1$, considering $1$ as the best result possible and 00 as the worst. The score for the parameters is computed in the same way on the Params column. The average of the scores obtained for 30- and 60-minute horizons are then used to obtain the final rank independently from the horizon considered.
TABLE III: Sequential Transformer performance on OhioT1DM and DiaTrend datasets in different training settings*
@@
as observable in Table I, the best rank is achieved by the proposed model with a single-modality approach, while the multimodal version achieves the third rank. This means that, on the OhioT1DM dataset, the Sequential Transformer proves to be optimal for real-life situations requiring edge computing, thanks to its extremely low dimensionality in terms of parameters and its performance considering only CGM. The second best performing model is Bi-GRU [29]. However, it is important to notice that even if the model is relatively small in terms of parameters and compared with other architectures, its dimension can still be a problem on wearable devices. Another fundamental analysis regards the identification of future adverse events. The proposed strategy with the Sequential Transformer can always achieve the best or among the best results in detecting hyper and hypo events. On this particular issue, it is possible to observe that the multimodal Sequential Transformer outperforms the Bi-GRU, thus being more suitable in a real-life scenario.
@@
Table II shows a trend similar to the OhioT1DM dataset. However, in this case, the Sequential Transformer takes both the best and second-best positions on the ranking. The single-modality version achieves the best ranking, with the multimodal version reaching the second place. As before, the Sequential Transformer in every version achieves the best or among the best performance in detecting adverse events. The only exception is TimesNet [27], which achieves the second-best performance in Hypo Sen. Nonetheless, its number of parameters makes it unsuitable for real-life scenarios where edge computing results in a better solution.
V-A Ablation studies
@@
The first rows of Table III report the scores obtained on the OhioT1DM dataset with different settings. The best-ranked method is represented by our Sequential Transformer using the proposed balancing strategy, implementing data augmentation, and, as determined before, using only CGM. The second-best is the multimodal version, with the same setting in terms of loss and data augmentation. It is interesting to notice that, based on the ranking, the balanced versions represent the first four best choices, demonstrating the effectiveness of the defined training strategy. Similar conclusions can be drawn for the DiaTrend dataset. The first rank is still taken by the single-modality setting with a balanced strategy and data augmentation. The second is the same setting but with a multimodal approach. Even in this case, it is possible to notice that the four best-ranked settings are all based on the balance training strategy, remarking once again on the importance of such implementation.
@@
V-B Clarke Error Grid Analysis
@@
V-C Deployment on wearable devices